OpenStudy (anonymous):

solve log10^(x+4) + log10 ^(x-2) = 1 I got x= -1 + sq root of 19 or -1 + sq root of 19. Just wanted to make sure I'm doing these right! thanks so much!

OpenStudy (lgbasallote):

how does this look like? \[\LARGE \log 10^{x+4} + \log 10^{x-2} = 1\]

OpenStudy (lgbasallote):

or is it \[\LARGE \log_{10} (x+4) + \log_{10} (x-2) = 1\]

OpenStudy (anonymous):

Yes, like the second equation, log base 10 :)

OpenStudy (anonymous):

\[log_{10}((x+4)(x-2))=1\]

OpenStudy (anonymous):

then use your log properties to change into exponential form

OpenStudy (lgbasallote):

i see..first you incorporate the product rule then \[\log_{10} (x+4)(x-2) = 1\] then you transform it into exponential form \[10^1 = (x+4)(x-2)\] \[10 = (x+4)(x-2)\]

OpenStudy (lgbasallote):

then you expand the right side

OpenStudy (anonymous):

dont forgot the domain x+4>0 and x-2>0 so ur domain is x>-4

OpenStudy (anonymous):

ah I see thanks!!, so would you put the answer using the quadratic formula??

OpenStudy (lgbasallote):

i think it's non-factorable so yeah

OpenStudy (anonymous):

Great thank you!!

OpenStudy (lgbasallote):

<tips imaginary hat>

OpenStudy (anonymous):

x= -1 - sq root of 19 and -1 + sq root of 19 but x=-1- sq root of 19 is not in domain so only answer is x=-1 + sq root of 19

OpenStudy (anonymous):

haha thank you both of you!! I would def give you both tips, however, I'm an unemployed pre-med student for now :)

OpenStudy (anonymous):

welcome :)